A 16 kg tuna fish moving horizontally to the right at 6 m/s swallows a 3 kg dalagang bukid that is swimming to the left at 8.5 m/s. What is the speed of the tuna fish immediately after, if the forces exerted on the fish by the water can be neglected?
1
Expert's answer
2020-11-23T05:28:37-0500

The total momentum of two fishes befor the impact was:


"p = m_1v_1 - m_2v_2"


where "m_1 = 16.6kg" is the mass of the tuna, "v_1 = 6m\/s" is the speed of the tuna, "m_2 = 3kg" is the mass of the dalagang bukid and "v_2 = 8.5m\/s" is the speed of the dalagang bukid.

The momentum of two fishes after the impact is:


"p' = (m_1 + m_2)v'"

where "v'" is their speed (tuna's speed, in fact) after the impact.

Let's apply the momentum conservation law:


"p = p'\\
m_1v_1 - m_2v_2 = (m_1 + m_2)v'\\
v' = \dfrac{m_1v_1 - m_2v_2}{m_1 + m_2}\\
\space \\
v' = \dfrac{16.6\cdot 6 - 3\cdot 8.5}{16.6 + 8.5} \approx 2.95m/s"

Answer. 2.95 m/s.